# The Spectrum of Gravitational Waves from Annihilating Domain Walls

### v1

@@ Abstract
Networks of cosmic domain walls can form in the early Universe as a consequence of the spontaneous breaking of discrete symmetries. We study the production of a cosmological background of gravitational waves (GWs) from such networks, when they annihilate due to a small explicit symmetry breaking term. Averaging over several 3+1-dimensional high-resolution lattice field simulations, we obtain a GW spectrum with the following characteristics: (1) a broad asymmetric peak, roughly located at frequency $f\sim 2H_{\rm gw}$, where $H_{\rm gw}$ is the Hubble rate at the end of GW production, shortly after annihilation, (2) a doubly broken power spectrum $\propto k^{-n}$, with initial slope $n\sim 0.5$ after the main peak and $n\sim 1.8$ at high $f$, while the low frequency region $f<f_{p}$ agrees with the causality behavior $\sim k^{3}$. Additionally, extending previous results, we find that GW production continues to be efficient until a value of the Hubble scale $H_{\text{g}w}$ that is roughly an order of magnitude smaller than the naive estimate $\sigma H=\Delta V$, where $\sigma$ is the wall tension and $\Delta V$ the size of the symmetry breaking term, thereby leading to a $O(100)$ larger GW signal. We find such results to be robust when changing the shape of the scalar field potential or including a time-dependent symmetry breaking term. Our findings have important implications for GW searches, especially in light of the reported evidence for a stochastic GW background in Pulsar Timing Array data.
cosmology, early Universe, non-perturbative effects, reheating, preheating, inflation
@@ I Introduction
The two-dimensional case, i.e. domain walls (DWs), is realized when a discrete symmetry is spontaneously broken in the cosmological evolution, and arguably provides the best example of such an opportunity. Indeed, DWs stand out compared to defects of lower dimensionality because of two reasons. First, they arise already in the simple field theory of a real scalar field with a $\mathbb{Z}_{2}$-symmetric potential (in contrast to monopoles and strings, which require more degrees of freedom). Second, while their evolution achieves a scaling regime which closely resembles that of cosmic strings [1, 2], whereby an approximately fixed number of Hubble-sized defects is present at each epoch in the cosmological evolution, their higher spatial dimensionality implies that their energy density redshifts slower than radiation and matter, i.e. $\rho_{\text{dw}}\sim\sigma H\sim a^{-2}$, where $\sigma$ is the domain wall mass per unit surface (tension), $H$ the Hubble expansion rate and $a$ is the scale factor. Such a slow dilution rules out the possibility of spontaneously broken, but otherwise exact, discrete symmetries in the radiation epoch [1], if the resulting DW tension is larger than about the $\text{MeV}^{3}$ scale. Nonetheless, as long as a sufficiently large explicit breaking of the discrete symmetry is present [5], that may be expected on general grounds, domain walls can decay in the early Universe, while they are still subdominant compared to the radiation background (see e.g. [9, 10, 11, 12, 13] for other possibilities to evade the so-called domain wall problem).
@@
Due to the non-linearities that are intrinsic to topological defects, a detailed characterization of the GW signal from domain walls requires numerical (lattice classical field theory) simulations of domain wall formation, evolution and annihilation in the expanding Universe [28]. A similar necessity has long been realized and addressed for other primordial GW sources, in particular first order phase transitions and local cosmic strings, see e.g. [29] for a recent collection of results. The first simulations with this aim were performed more than a decade ago [17], also for hybrid string-wall networks [30], focusing on the scaling regime (see also [31] for string-wall network simulations in the context of dark matter). The results of these works have since then been used both for GW searches (in LVK [32] and PTA data [26, 27]) as well as for phenomenological analyses of particle physics models. Very recently, updated numerical studies have appeared, focused on: specific axion models for DW annihilation [33]; dark matter production from the decay of domain walls , [34]; the estimation of the fraction of Primordial Black Holes (PBHs) that can be produced by DW collapse [19]; the GW signal from the scaling regime [20], i.e. not including the annihilation phase, and the GW signal from melting domain walls [35].
@@ III Main model: $\lambda\phi^{4}$ with T-independent bias
where $q\ll 1$ is a time-independent constant.We use a cubic, rather than linear, symmetry breaking potential to ensure that no bias is introduced in the initial population of the two minima. [^2] Alternatively, one can choose a linear bias and set the initial homogeneous value of the field to coincide with the maximum of $V$, rather than setting $\phi_{i}=0$ as we do. We have checked that this alternative choice does not affect our results.. The potential (12) is shown by the blue curve in Fig. 1, for the actual value of $q$ used in our simulations. In the presence of $V_{\text{bias}}$, the energies of the two minima differ by $\Delta V=2q(1+9q^{2})^{3/2}v^{4}$. It is useful to define a quantity $\eta_{\Delta V}$ as the (conformal) time at which the vacuum pressure equals the energy density of one Hubble sized wall, i.e. $\sigma H(\eta_{\Delta V})=\Delta V$. Roughly speaking, we expect the network to start deviating from the scaling regime around this time. This is confirmed by inspection of the area parameter, as shown in the left panel of Fig. 2, where we have chosen $\eta_{\Delta V}=14$. The area parameter in the biased case coincides with its value in the unbiased case roughly until time $\eta_{\Delta V}$, then decreases as the network undergoes annihilation.
@@
[…]
@@
Figure 6: Red dots show the amplitude of the averaged GW spectrum at different spectral bins for the biased $\phi^{4}$ potential, extracted at time $\eta_{f}=47$, as a function of the wavenumber $x\equiv k/(2\pi aH)$. The bars indicate the error coming from averaging seven different realizations. The blue/purple dashed lines show the best fit to the double/single broken power law templates, given by Eqs. (15) and (14), respectively.
@@
where $\alpha$ and $\beta$ capture the spectral slopes at $x\ll x_{p}$ and $x\gg x_{p}$ respectively, and $\delta$ is the width around the maximum. We set $\alpha=3$ because of causality. Overall, from the fit to the data we can extract the value of three parameters describing the shape of the spectrum, i.e. $x_{p}$, $\beta$ and $\delta$, as well as the one additional parameter to describe its amplitude, the efficiency parameter $\epsilon$. We obtain the following constraints:
@@
Biased $\phi^{4}$ potential: double bpl template (15)
@@
Table 2: Posteriors for the parameters of the double broken power law template (15) for the biased $\phi^{4}$ potential, that provides an excellent fit to our numerical spectra.
In Fig. 6 we compare the results of both fits with the averaged GW spectrum extracted from the lattice. It can be appreciated that the peaked double broken power law template provides a significantly better fit to the numerical data than the simpler peaked template without a breaking at high frequencies. We then focus on the posteriors for the fitting parameters obtained with this spectral shape.
@@
    First, the GW signal from annihilating walls peaks at a wavenumber that is significantly larger than in the unbiased case. In particular, we find $x_{p}\simeq 2.2$, whereas in the scaling regime we find $x_{p}\simeq 1$, see App. A. This implies that the peak frequency of the GW signal today is roughly twice as large as the value that has been used so far in searches for GWs from domain walls in LVK and PTA datasets [32, 26, 27]. We notice that this roughly agrees with the result presented in [33], which performed a single realization with a temperature-dependent linear bias to annihilate the network. In App. B.5 we compare the results of [33] with ours, while in Sec. IV we consider the case of a temperature-dependent bias.
@@
Before moving on to considering other DW models, let us discuss potential pitfalls of our approach and the associated potential impact on our results. For reasons of computational capabilities, we are not able to extensively test the validity of our results for different bias sizes (i.e. different values of $\eta_{\Delta V}$). One interesting and relevant question that requires further investigation is then whether the parameter $x_{b}$ actually depends on $\eta_{\Delta V}$. While we are not able to confidently address this question in this work, we notice that $x_{b}/x_{p}$ is close to the ratio between the time at which GW production starts deviating significantly from the scaling behavior (around $\eta\approx 28$, see Fig. 5) and the time at which production ends. We expect this ratio to remain roughly constant as we change $\eta_{\Delta V}$ (we have checked that this is the case with one other choice of bias size, see App. B.4, see also [19]). Therefore, if the existence of the breaking point is related to the separation between $\eta_{\Delta V}$ and the final time of GW production, we expect $x_{b}$ to be roughly independent from $\eta_{\Delta V}$. The ratio $x_{b}/x_{p}$ is particularly relevant for observability at multiple GW observatories: if the main peak is indeed observed at e.g. PTAs, the UV tail would be observable at LISA if $x_{b}\gg x_{p}$, i.e. the intermediate slope extends for several orders of magnitude with $\beta\lesssim 1$. If instead the signal features a transition to a steeper slope at $x_{b}\simeq 2.8\leavevmode\nobreak\ x_{p}$ as we see in our simulations, then the UV tail would not be observable at LISA.
@@ IV.1 Cosine potential
whose periodicity is twice as large as that of potential (16). The symmetry breaking potential is out of phase with respect to (16), in such a way that the the minimum at $\phi=-v$ is unaltered, while the minimum at $\phi=+v$ is lifted by $\Delta V\simeq 2p\lambda v^{4}$. The total potential is shown in Fig. 1 by the orange curve. The conformal time $\eta_{\Delta V}$ is then determined by setting $\sigma H=\Delta V$ as usual. The potential (17) induces a shift in the position of the maximum of the potential, differently from the $\phi^{4}$ case with the cubic bias. In order to avoid a bias in the initial population of the two minima, we thus set initial conditions for $\phi$ such that its homogeneous initial value coincides with the maximum of the total potential $V=V+V_{\text{bias}}$, rather than setting $\phi=0$ as in Sec. III.
@@
Table 3: Posteriors for the parameters of the double broken power law template (15) in the biased cosine case.
@@ V Conclusions and Observational Impact
$\Omega_{\text{gw}}(f_{p})\simeq 10^{-10}\epsilon\left(\frac{g_{*}(T_{\text{gw}})}{10.75}\right)^{-1/3}\left(\frac{\alpha_{\text{gw}}}{0.01}\right)^{2}.$ (22)
@@
Additionally, we were able to improve on the determination of $T_{\text{gw}}$, i.e. the temperature at which efficient GW production stops (and thus the temperature that sets the position of the peak frequency of the relic GW spectrum (V) as well as the amplitude of $\alpha_{\star}$), in terms of the DW tension $\sigma$ and the vacuum energy difference between the two minima $\Delta V$. Previous work [19] has already found that $T_{\text{gw}}<T_{\Delta V}$, where $T_{\Delta V}$ is the temperature defined by the commonly employed condition $\sigma H=\Delta V$ for DW annihilation. In our simulations with $\eta_{\Delta V}=14$ we find $\eta_{\text{gw}}=44.4\pm 1.0$, i.e. $\eta_{\text{gw}}\simeq 3.2\eta_{\Delta V}$. Therefore, we find an even larger delay in GW production with respect to what was already noticed in [19], attributing the differences to the improvements in our simulations. A careful extension of this result to a general bias size requires further studies, although we expect the amount of the delay to remain constant. Under the assumptions that the delay remains constant when changing the size of the bias term (see [19] and our App. B.4), we deduce
@@
Note added: very recently, a preprint [45] has also presented results on gravitational waves from annihilating domain walls, based on lattice simulations of $N=2048$ points per dimension.

### v3

@@ Abstract
Networks of cosmic domain walls can form in the early Universe as a consequence of the spontaneous breaking of discrete symmetries. We study the production of a cosmological background of gravitational waves (GWs) from such networks, when they annihilate due to a small explicit symmetry breaking term. Averaging over several 3+1-dimensional high-resolution lattice field simulations, we obtain a GW spectrum with the following characteristics: (1) a broad asymmetric peak, roughly located at frequency (at the time of emission) $f\sim 2H_{\rm gw}$, where $H_{\rm gw}$ is the Hubble rate at the end of GW production, shortly after annihilation, (2) a doubly broken power law spectrum $\propto k^{-n}$, with initial slope $n\sim 0.5$ after the main peak and $n\sim 1.8$ at high $f$, while the low frequency region $f<f_{p}$ agrees with the causality behavior $\sim k^{3}$. Additionally, extending previous results, we find that GW production continues to be efficient until a value of the Hubble scale $H_{\text{g}w}$ that is roughly an order of magnitude smaller than the naive estimate $\sigma H=\Delta V$, where $\sigma$ is the wall tension and $\Delta V$ the size of the symmetry breaking term, thereby leading to a $O(100)$ larger GW signal. We find such results to be robust when changing the shape of the scalar field potential or including a time-dependent symmetry breaking term. Our findings have important implications for GW searches, especially in light of the reported evidence for a stochastic GW background in Pulsar Timing Array data.
@@ I Introduction
The two-dimensional case, i.e. domain walls (DWs), is realized when a discrete symmetry is spontaneously broken in the cosmological evolution, and arguably provides the best example of such an opportunity. Indeed, DWs stand out compared to defects of lower dimensionality because of two reasons. First, they arise already in the simple field theory of a real scalar field with a $\mathbb{Z}_{2}$-symmetric potential (in contrast to monopoles and strings, which require more degrees of freedom). Second, while their evolution achieves a scaling regime which closely resembles that of cosmic strings [1, 2], whereby an approximately fixed number of Hubble-sized defects is present at each epoch in the cosmological evolution, their higher spatial dimensionality implies that their energy density redshifts slower than radiation and matter, i.e. $\rho_{\text{dw}}\sim\sigma H\sim a^{-2}$, where $\sigma$ is the domain wall mass per unit surface (tension), $H$ the Hubble expansion rate and $a$ is the scale factor. Such a slow dilution rules out the possibility of spontaneously broken, but otherwise exact, discrete symmetries in the radiation epoch [1], if the resulting DW tension is larger than about the $\text{MeV}^{3}$ scale. Nonetheless, as long as a sufficiently large explicit breaking of the discrete symmetry is present [5], that may be expected on general grounds, domain walls can decay in the early Universe, while they are still subdominant compared to the radiation background (see e.g. [9, 10, 11, 12, 13, 14, 15] for other possibilities to evade the so-called domain wall problem).
@@
Due to the non-linearities that are intrinsic to topological defects, a detailed characterization of the GW signal from domain walls requires numerical (lattice classical field theory [30]) simulations of domain wall formation, evolution and annihilation in the expanding Universe. A similar necessity has long been realized and addressed for other primordial GW sources, in particular first order phase transitions and local cosmic strings, see e.g. [31] for a recent collection of results. The first simulations with this aim were performed more than a decade ago [19], also for hybrid string-wall networks [32], focusing on the scaling regime (see also [33, 34, 35] for DW simulations without GWs). The results of these works have since then been used both for GW searches (in LVK [36] and PTA data [28, 29]) as well as for phenomenological analyses of particle physics models. Very recently, updated numerical studies have appeared, focused on: specific axion models for DW annihilation [37]; dark matter production from the decay of domain walls [38]; the estimation of the fraction of Primordial Black Holes (PBHs) that can be produced by DW collapse [21] (see also [39]); the GW signal from the scaling regime, both in matter [40] and radiation-domination [41, 22], i.e. not including the annihilation phase, and the GW signal from melting domain walls [42].
@@ III Main model: $\lambda\phi^{4}$ with T-independent bias
where $q\ll 1$ is a time-independent constant. We use a cubic, rather than linear, symmetry breaking potential to ensure that no bias is introduced in the initial population of the two minima. [^2] Alternatively, one can choose a linear bias and set the initial homogeneous value of the field to coincide with the maximum of $V$, rather than setting $\phi_{i}=0$ as we do. We have checked that this alternative choice does not affect our results.. The potential (12) is shown by the blue curve in Fig. 1, for the actual value of $q$ used in our simulations. In the presence of $V_{\text{bias}}$, the energies of the two minima differ by $\Delta V=qv^{4}(4\lambda^{2}+9q^{2})^{3/2}/(4\lambda^{3})$. It is useful to define a quantity $\eta_{\Delta V}$ as the (conformal) time at which the vacuum pressure equals the energy density of one Hubble sized wall, i.e. $\sigma H(\eta_{\Delta V})=\Delta V$. Roughly speaking, we expect the network to start deviating from the scaling regime around this time. This is confirmed by inspection of the area parameter, as shown in the left panel of Fig. 2, where we have chosen $\eta_{\Delta V}=14$. The area parameter in the biased case coincides with its value in the unbiased case roughly until time $\eta_{\Delta V}$, then decreases as the network undergoes annihilation.
@@
Anticipating slightly the results of our simulations, we find that, for $\eta_{\Delta V}=14$, GW production continues until very late times, around $\eta\approx 45$. According to the conditions in (10)-(11), by requiring $c(\eta_{\text{max}})\gtrsim 1$, and fixing $L\gtrsim 60$ in order to have sufficient IR resolution, these times are barely achievable with the box sizes that we are able to simulate. We are thus constrained to choose $\eta_{\Delta V}=14$ (we have nonetheless obtained results for the scalar spectrum with several other choices of bias size, see App. B.5, that support the conclusion that the shape of the GW power spectrum should not be strongly affected by the bias size).
Nonetheless, we find that most of the walls disappear by the time $\eta\approx 36$, as can be appreciated in Fig. 3, where we present snapshots of the gradient energy distribution at different times across a two-dimensional slice of the three-dimensional lattice. Correspondingly, we have indeed checked that the condition that the DWs width must be resolved until the end of the simulation is indeed overconservative: namely, comparing the results of a simulation with $L=80$ (where DWs are resolved until $\eta_{\rm max}=32$) to those with $L=60$ (where DWs are resolved until $\eta_{\rm max}=47$) we notice only marginal differences in our results. Therefore, we obtain most of our results with the parameters $N=3060,L=80$, which importantly provide a better IR coverage for the GW spectrum until late times. We do nonetheless test our results against individual realizations with smaller $L$, see App. B.4 for more details.
@@
Figure 6: Red dots show the amplitude of the averaged GW spectrum at different spectral bins for the biased $\phi^{4}$ potential, extracted at time $\eta_{f}=47$, as a function of the wavenumber $x\equiv k/(2\pi aH)$. The bars indicate the error coming from averaging seven different realizations. The blue/purple dashed lines show the best fit to the doubly/single broken power law templates, given by Eqs. (15) and (14), respectively.
@@
where $\alpha$ and $\beta$ capture the spectral slopes at $x\ll x_{p}$ and $x\gg x_{p}$ respectively, and $\delta$ is the width around the maximum. We set $\alpha=3$ because of causality [50, 51]. Overall, from the fit to the data we can extract the value of three parameters describing the shape of the spectrum, i.e. $x_{p}$, $\beta$ and $\delta$, as well as the one additional parameter to describe its amplitude, the efficiency parameter $\epsilon$. We obtain the following constraints:
@@
Biased $\phi^{4}$ potential: doubly bpl template (15)
@@
Table 2: Posteriors for the parameters of the doubly broken power law template (15) for the biased $\phi^{4}$ potential, that provides an excellent fit to our numerical spectra.
In Fig. 6 we compare the results of both fits with the averaged GW spectrum extracted from the lattice. It can be appreciated that the peaked doubly broken power law template provides a significantly better fit to the numerical data than the simpler peaked template without a breaking at high frequencies. We then focus on the posteriors for the fitting parameters obtained with this spectral shape.
@@
    First, the GW signal from annihilating walls peaks at a wavenumber that is significantly larger than in the unbiased case. In particular, we find $x_{p}\simeq 2.2$, whereas in the scaling regime we find $x_{p}\simeq 1.2$, see App. A. This implies that the peak frequency of the GW signal today is roughly twice as large as the value that has been used so far in searches for GWs from domain walls in LVK and PTA datasets [36, 28, 29]. We notice that this roughly agrees with the result presented in [37], which performed a single realization with a temperature-dependent linear bias to annihilate the network. In App. B.6 we compare the results of [37] with ours, while in Sec. IV we consider the case of a temperature-dependent bias.
@@
Before moving on to considering other DW models, let us discuss potential pitfalls of our approach and the associated potential impact on our results. For reasons of computational capabilities, we are not able to extensively test the validity of our results for different bias sizes (i.e. different values of $\eta_{\Delta V}$). One interesting and relevant question that requires further investigation is then whether the parameter $x_{b}$ actually depends on $\eta_{\Delta V}$. While we are not able to confidently address this question in this work, we notice that $x_{b}/x_{p}$ is close to the ratio between the time at which GW production starts deviating significantly from the scaling behavior (around $\eta\approx 28$, see Fig. 5) and the time at which production ends. We expect this ratio to remain roughly constant as we change $\eta_{\Delta V}$ (we have checked that this is the case with one other choice of bias size, and we have investigated the effects of different bias sizes on the scalar field spectrum, see App. B.5, see also [21]). Therefore, if the existence of the breaking point is related to the separation between $\eta_{\Delta V}$ and the final time of GW production, we expect $x_{b}$ to be roughly independent from $\eta_{\Delta V}$. The ratio $x_{b}/x_{p}$ is particularly relevant for observability at multiple GW observatories: if the main peak is indeed observed at e.g. PTAs, the UV tail would be observable at LISA if $x_{b}\gg x_{p}$, i.e. the intermediate slope extends for several orders of magnitude with $\beta\lesssim 1$. If instead the signal features a transition to a steeper slope at $x_{b}\simeq 2.8\leavevmode\nobreak\ x_{p}$ as we see in our simulations, then the UV tail would not be observable at LISA.
@@ IV.1 Cosine potential
whose periodicity is twice as large as that of potential (16). The symmetry breaking potential is out of phase with respect to (16), in such a way that the minimum at $\phi=-v$ is unaltered, while the minimum at $\phi=+v$ is lifted by $\Delta V\simeq 2p\lambda v^{4}$. The total potential is shown in Fig. 1 by the orange curve. The conformal time $\eta_{\Delta V}$ is then determined by setting $\sigma H=\Delta V$ as usual. The potential (17) induces a shift in the position of the maximum of the potential, differently from the $\phi^{4}$ case with the cubic bias. In order to avoid a bias in the initial population of the two minima, we thus set initial conditions for $\phi$ such that its homogeneous initial value coincides with the maximum of the total potential $V=V+V_{\text{bias}}$, rather than setting $\phi=0$ as in Sec. III.
@@
Table 3: Posteriors for the parameters of the doubly broken power law template (15) in the biased cosine case.
@@ V Conclusions and Observational Impact
$\Omega_{\text{gw}}(f_{p})h^{2}\simeq 10^{-10}\epsilon\left(\frac{g_{*}(T_{\text{gw}})}{10.75}\right)^{-1/3}\left(\frac{\alpha_{\text{gw}}}{0.01}\right)^{2}.$ (22)
@@
Additionally, we were able to improve on the determination of $T_{\text{gw}}$, i.e. the temperature at which efficient GW production stops (and thus the temperature that sets the position of the peak frequency of the relic GW spectrum (V) as well as the amplitude of $\alpha_{\star}$), in terms of the DW tension $\sigma$ and the vacuum energy difference between the two minima $\Delta V$. Previous work [21] has already found that $T_{\text{gw}}<T_{\Delta V}$, where $T_{\Delta V}$ is the temperature defined by the commonly employed condition $\sigma H=\Delta V$ for DW annihilation. In our simulations with $\eta_{\Delta V}=14$ we find $\eta_{\text{gw}}=44.4\pm 1.0$, i.e. $\eta_{\text{gw}}\simeq 3.2\eta_{\Delta V}$. Therefore, we find an even larger delay in GW production with respect to what was already noticed in [21], attributing the differences to the improvements in our simulations. A careful extension of this result to a general bias size requires further studies, although we expect the amount of the delay to remain constant. Under the assumptions that the delay remains constant when changing the size of the bias term (see [21] and our App. B.5), we deduce [^7] In [21], the parameter $\eta_{\text{ann}}$ was introduced, corresponding to the time at which the fraction of the simulation box in the false vacuum (i.e. in $\phi>0$) starts decreasing exponentially. It was found that $\eta_{\text{ann}}\gtrsim\eta_{\Delta V}$. One would expect $\eta_{\text{gw}}/\eta_{\text{ann}}$ to be independent of $\Delta V$, rather than $\eta_{\text{gw}}/\eta_{\Delta V}$. Nonetheless, if $\eta_{\text{ann}}$ and $\eta_{\Delta V}\propto\Delta V^{1/2}$ have the same dependence on $\Delta V$, then the extrapolation (23) is justified. By fitting the results presented in Table 1 of [21], we find that $\eta_{\rm ann}\propto(\Delta V)^{-(0.40\pm 0.05)}$. This result is compatible up to two standard deviations with the standard expectation, although it is also compatible with the recent claim by [55] $\eta_{\rm ann}\propto(\Delta V)^{1/3}$ (that appeared after our work, and is obtained with lower resolution than [21]). While a dedicated analysis is required to reach a conclusion, we notice that a different scaling would only affect the relation between the amplitude of the GW signal and the parameters $\sigma$ and $\Delta V$. The shape of the spectrum would not be altered.
@@
Note added: shortly before our paper appeared on the arXiv, the preprint [56] has also presented results on gravitational waves from annihilating domain walls, based on lattice simulations of $N=2048$ points per dimension.